# Selective Vision is the Challenge for Visual Reasoning: A Benchmark for Visual Argument Understanding

### v1

@@ Abstract
We collect and release VisArgs, an annotated corpus designed to make explicit the (usually implicit) structures underlying visual arguments. VisArgs includes 1,611 images accompanied by three types of textual annotations: 5,112 visual premises (with region annotations), 5,574 commonsense premises, and reasoning trees connecting them to a broader argument. We propose three tasks over VisArgs to probe machine capacity for visual argument understanding: localization of premises, identification of premises, and deduction of conclusions. Experiments demonstrate that 1) machines cannot fully identify the relevant visual cues. The top-performing model, GPT-4-O, achieved an accuracy of only 78.5%, whereas humans reached 98.0%. All models showed a performance drop, with an average decrease in accuracy of 19.5%, when the comparison set was changed from objects outside the image to irrelevant objects within the image. Furthermore, 2) this limitation is the greatest factor impacting their performance in understanding visual arguments. Most models improved the most when given relevant visual premises as additional inputs, compared to other inputs, for deducing the conclusion of the visual argument. Our code and data are publicly released at https://github.com/JiwanChung/VisArgs.
@@ 2 Related Work
Visual arguments are arguments built on visual medium Boland (2005). Unlike typical images, a visual argument is intentionally organized to persuade viewers to a certain conclusion Birdsell and Groarke (1996); Boland (2005). This work builds upon to ongoing debates in the human studies literature about the nature of visual arguments Johnson (2003); Tseronis (2018). Our results (§ 5) suggest that understanding visual arguments requires focusing on a subset of the visual context: not all visual cues contribute, and identifying the relevant ones is the key necessity. This task one of selective vision: the human capability to focus on behaviorally relevant stimuli. Desimone and Duncan (1995). Examples of visual arguments are prevalent in advertisements Kjeldsen (2012); Zhang et al. (2018); Ye et al. (2019), cartoons Birdsell and Groarke (2007), mathematical educations Inglis and Mejía-Ramos (2009), and, arguably, diagrams Kembhavi et al. (2016); Alikhani and Stone (2018).
@@ 3.1 Annotation Process
Collecting Images. We manually collect around 1,600 images from Pinterest.222www.pinterest.com Starting with keyword-based searches (e.g. creative ads), we expanded our collection by exploring related images. Cartoons (which often contain visual arguments Birdsell and Groarke (1996)) were sourced from a dedicated website.333www.cartoonmovement.com We manually collected around 1,600 cartoons from various categories, including politics, education, and environment. For both categories, we followed previous work Schuhmann et al. (2022); Lee et al. (2021) by including URLs to the images to comply with licensing terms. Refer to Appendix A for details.
Describing Visual Premises. The next step is to explicitly describe the visual argument within each image. However, during the early stages of our annotation process, we discovered that although humans can naturally understand visual arguments, they often find it challenging to articulate their interpretation into structured argumentation trees. Therefore, we used an AI model (GPT-4-O) to generate initial candidates. Human workers then select and modify these initial annotations, as shown in Fig. 3. To facilitate this process, we break down the annotation into two steps: describing the visual premises and specifying the argument structure.
@@ 3.2 Data Analysis
Visual Cues vs. Dense Captioning. In theory, selective attention to visual premises could be collapsed into an NLP problem by describing everything in an image. To test this counter-hypothesis, we manually check how often the visual premises are contained in the outputs of detailed captioning models. We include three baselines here: a generalist (LLaVA-Next Liu et al. (2024b)), a specialist (ShareCaptioner Chen et al. (2023)), and LLaVA-LLaMa3 XTuner Contributors (2023) fine-tuned on a detailed captioning corpus (DOCCI Onoe et al. (2024))444huggingface.co/gokaygokay/llava-llama3-docci. Tab. 1 summarizes our manual inspection of 100 images, showing that the detailed captions insufficiently capture the visual premises, with the hit rate staying below $15\%$ for all models.
Safety. Since we did not initially filter for safety, we now analyze the safety of VisArgs using standard models. For textual safety, we utilize the Perspective API555www.perspectiveapi.com; June 2024 version., and for visual domains, we employ LAION-Safety666www.github.com/LAION-AI/LAION-SAFETY. The toxicity scores for textual descriptions were $0.03$ for visual premises and $0.07$ for conclusions. Also, given the threshold of $0.7$, no descriptions and visual premises were classified as toxic. Furthermore, only $71$ among $1611$ images are classified as unsafe. Manual inspection reveals that such “unsafe" images were social campaigns advocating against the harmful behaviors which presumably triggered the LAION detector.
@@ 5.1 Localization of Premises
Metrics and Models. For closed-set grounding, which is an N-way classification task, the goal is to match the given description with the correct bounding box. To evaluate standard image-text matching algorithms (e.g. CLIP), we crop the regions accordingly. The models for this task include various CLIP-based models (CLIP Radford et al. (2021) with different backbones and SigLIP Zhai et al. (2023)) and a multitask model OFA Wang et al. (2022). For open-set grounding, which is to locate an object in an image based on a natural language expression, we instruct the models to output bounding box coordinates and we compare them to the ground truth region. A predicted coordinate is considered correct if its intersection over union with the gold label is at least (IoU $\geq 0.5$). We use a diverse set of models that support local region output formats, UNINEXT-H Yan et al. (2023), LISA Lai et al. (2023), Unified-IO-2 Lu et al. (2023), OFA, MM-G-DINO Liu et al. (2023b).
@@ Appendix B Analyzing Topic Diversity
Labelling Data. We use a pretrained language model to classify visual premises (VP) and conclusions (C) in a zero-shot manner. We provide the following input to the LLaMA-3777meta-llama/Meta-Llama-3-8B-Instruct LLM:

### v3

@@ Abstract
We present VisArgs  [^1] Data: https://huggingface.co/datasets/jiwan-chung/visargs, a dataset of 1,611 images annotated with 5,112 visual premises (with regions), 5,574 commonsense premises, and reasoning trees connecting them into structured arguments. We propose three tasks for evaluating visual argument understanding: premise localization, premise identification, and conclusion deduction. Experiments [^2] Code: https://github.com/JiwanChung/VisArgs show that 1) machines struggle to capture visual cues: GPT-4-O achieved 78.5% accuracy, while humans reached 98.0%. Models also performed 19.5% worse when distinguishing between irrelevant objects within the image compared to external objects. 2) Providing relevant visual premises improved model performance significantly.
@@ 2 Related Work
Visual arguments are arguments built on visual medium Boland (2005). Unlike typical images, a visual argument is intentionally organized to persuade viewers to a certain conclusion Birdsell and Groarke (1996); Boland (2005). This work builds upon to ongoing debates in the human studies literature about the nature of visual arguments Johnson (2003); Tseronis (2018). Our results (§ 5) suggest that understanding visual arguments requires focusing on a subset of the visual context: not all visual cues contribute, and identifying the relevant ones is the key necessity. This task one of selective vision: the human capability to focus on behaviorally relevant stimuli. Desimone and Duncan (1995). Examples of visual arguments are prevalent in advertisements Kjeldsen (2012); Zhang et al. (2018); Hussain et al. (2017); Ye et al. (2019), cartoons Birdsell and Groarke (2007), mathematical educations Inglis and Mejía-Ramos (2009), and, arguably, diagrams Kembhavi et al. (2016); Alikhani and Stone (2018). Liu et al. Liu et al. (2022) also investigate arguments conveyed through images. However, our work focuses on visual argument, requiring that the argumentative content be primarily communicated through visual elements rather than relying solely on accompanying captions or written text. Furthermore, we provide explicit annotations of the argumentation structure in a tree format, facilitating detailed, hierarchical analysis of the model’s ability to comprehend visual arguments step-by-step.
@@ 3.1 Annotation Process
Collecting Images. Our primary criterion was to select images that enable human annotators to easily and accurately interpret both the visual premise and the corresponding conclusion, thereby clarifying the argumentative structure within the image. Also, we ruled out samples with scene text within the images that directly describe the conclusion. We manually collect around 1,600 images following these criteria from Pinterest.444www.pinterest.com Starting with keyword-based searches (e.g. creative ads), we expanded our collection by exploring related images. Cartoons (which often contain visual arguments Birdsell and Groarke (1996)) were sourced from a dedicated website.555www.cartoonmovement.com We manually collected around 1,600 cartoons from various categories, including politics, education, and environment. We include URLs to the images to comply with licensing terms following previous work Schuhmann et al. (2022); Lee et al. (2021). Refer to Appendix A for details.
Describing Visual Premises. The next step is to explicitly describe the visual argument within each image. However, during the early stages of our annotation process, we discovered that although humans can naturally understand visual arguments, they often find it challenging to articulate their interpretation into structured argumentation trees. Therefore, we used an AI model (GPT-4-O) to generate initial candidates. Human workers then select and modify these initial annotations, as shown in Fig. 3. The human annotators could optionally incorporate new visual premises when necessary: $\sim 21\%$ of images had their set of visual premises expanded through this process. To facilitate this process, we break down the annotation into two steps: describing the visual premises and specifying the argument structure.
@@ 3.2 Data Analysis
Number of Samples Image Size (pixels)
Category Total W/ Text Width Height
Ads 914 389 877.2 969.7
Cartoons 697 218 480.0 427.6
Table 1: Overview of dataset statistics. W/ Text indicates the subset of images containing scene text.
@@
Visual Cues vs. Dense Captioning. In theory, selective attention to visual premises could be collapsed into an NLP problem by describing everything in an image. To test this counter-hypothesis, we manually check how often the visual premises are contained in the outputs of detailed captioning models. We include three baselines here: a generalist (LLaVA-Next Liu et al. (2024b)), a specialist (ShareCaptioner Chen et al. (2023)), and LLaVA-LLaMa3 XTuner Contributors (2023) fine-tuned on a detailed captioning corpus (DOCCI Onoe et al. (2024))666huggingface.co/gokaygokay/llava-llama3-docci. Tab. 2 summarizes our manual inspection of 100 images, showing that the detailed captions insufficiently capture the visual premises, with the hit rate staying below $15\%$ for all models.
Safety. Since we did not initially filter for safety, we now analyze the safety of VisArgs using standard models. For textual safety, we utilize the Perspective API777www.perspectiveapi.com; June 2024 version., and for visual domains, we employ LAION-Safety888www.github.com/LAION-AI/LAION-SAFETY. The toxicity scores for textual descriptions were $0.03$ for visual premises and $0.07$ for conclusions. Also, given the threshold of $0.7$, no descriptions and visual premises were classified as toxic. Furthermore, only $71$ among $1611$ images are classified as unsafe. Manual inspection reveals that such “unsafe" images were social campaigns advocating against the harmful behaviors which presumably triggered the LAION detector.
@@ 5.1 Localization of Premises
Metrics and Models. We define open-set evaluation as a setting in which models are required to generate bounding box coordinates without relying on a predefined candidate list. As a result, models used for open-set and closed-set evaluations are architecturally distinct, since models lacking an explicit generative head, such as CLIP Radford et al. (2021), are not compatible with open-set evaluation due to their dependence on a candidate region list for text-to-region matching.
For closed-set grounding, which is an N-way classification task, the goal is to match the given description with the correct bounding box. To evaluate standard image-text matching algorithms (e.g. CLIP), we crop the regions accordingly. The models for this task include various CLIP-based models (CLIP Radford et al. (2021) with different backbones and SigLIP Zhai et al. (2023)) and a multitask model OFA Wang et al. (2022). The CLIP-based models are adapted as follows: for each candidate object region (specified by bounding box coordinates), the corresponding regions are cropped from the image to create region-level image representations. Image features for each cropped region are then extracted using a CLIP-based encoder, while text features are obtained by encoding the input description using the same model. Cosine similarity is calculated between the text feature and each region-level image feature. The region with the highest similarity score is selected as the predicted match.
For open-set grounding, which is to locate an object in an image based on a natural language expression, we instruct the models to output bounding box coordinates and we compare them to the ground truth region. A predicted coordinate is considered correct if its intersection over union with the gold label is at least (IoU $\geq 0.5$). We use a diverse set of models that support local region output formats, UNINEXT-H Yan et al. (2023), LISA Lai et al. (2023), Unified-IO-2 Lu et al. (2023), OFA, MM-G-DINO Liu et al. (2023b).
@@ 5.4 Diagnostics
Reliance on OCR Capabilites. To examine how demanding VisArgs is on OCR capabilities, we employ a lightweight OCR detector Du et al. (2021) to detect bounding boxes of the visual premises without leveraging text annotations as input. Even this simplified model achieves $82.77\%$ accuracy in image-wise evaluation, where an image is considered correctly detected only if all visual premises within it are identified. Typical failure cases are illustrated in Fig. 6.
Figure 6: Left: OCR detection results. Right: Ground truth text instances missed by the model (highlighted in red). Most detection errors are attributed to Out-of-Domain cases such as calligraphy, handwritten text, or text that is too small for the model to detect, despite being distinguishable to the human eye.
@@ 6 Conclusion
We encourage future research to extend this work by exploring a wider range of visual arguments and incorporating more diverse cultural and linguistic contexts.
Acknowledgment
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University) and RS-2024-00353131) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00354218).
@@ Appendix A Data Annotation Details
Inter-Annotator Agreement. The dataset annotation was conducted by two primary human annotators, with a third evaluator assigned to assess annotation quality and consistency. To measure inter-annotator agreement, 100 samples were randomly selected and re-annotated by a secondary annotator different from the original. Subsequently, the third evaluator independently assessed the equivalence of each annotated sample.
Given that the annotations comprise premise sets and natural language conclusions, rather than numerical scores, traditional metrics such as Cohen’s kappa are not applicable. Instead, we measured inter-annotator agreement using two distinct criteria: equivalence of visual premise sets and equivalence of natural language conclusions. The Jaccard similarity index was employed to quantify the overlap between visual premise sets, while a binary equivalence test was used to evaluate alignment in the conclusions.
The results, as presented in Tab. 9, demonstrate a high degree of inter-annotator agreement. Qualitative analysis revealed that observed discrepancies primarily stemmed not from substantive semantic variations but from differences in how annotators segmented a single concept into one or more visual premises.
@@ Appendix B Analyzing Topic Diversity
Labelling Data. We use a pretrained language model to classify visual premises (VP) and conclusions (C) in a zero-shot manner. We provide the following input to the LLaMA-3999meta-llama/Meta-Llama-3-8B-Instruct LLM:
@@ Appendix C Experiment Details
Visual Premise Conclusion
Human-Human 0.78 0.96
Machine-Machine 0.51 0.88
Table 9: Inter-annotator agreement results based on human evaluation, quantified using the Jaccard Similarity Index for set-level comparisons.